OpenStudy (anonymous):

The region R in the first quandrant is enclosed by the lines x=0 and y=5 and the curve y=x^2+1. The volume of the solid generated when R is revolved about the y-axisi is a) 6pi b) 8pi c) 34pi/3 d) 16pi e) 544pi/15

OpenStudy (anonymous):

I am not understanding what im doing wrong but so far my awnser for some reason ends up being 4pi

OpenStudy (kainui):

Sure, so show me what you actually integrated or did and I can help you figure out where you went wrong. These are my favorite calculus questions!

OpenStudy (anonymous):

Well it being revoled around the y-axis I changed the formula to x=sqrt(y-1) so I intergrated this \[\pi \int\limits_{1}^{5} (\sqrt(y-1)^2dx\]

OpenStudy (anonymous):

Oh wait I just relized I forgot to take the anti derivative and I got 8pi

OpenStudy (anonymous):

Thats what I did wrong this whole time

OpenStudy (anonymous):

so then \[\pi([25/2 - 5] - [1/2 -1])=8\pi\]

OpenStudy (anonymous):

originally i did \[\pi([5-1]-[1-1])=4\pi\]

OpenStudy (anonymous):

Thanks anyways

OpenStudy (kainui):

Haha alright.